OpenStudy (anonymous):

Solve for y. y*(sqrt(1-y^2)) = 0.5

OpenStudy (anonymous):

This is the equation: \[y \sqrt{1-y ^{2}}= 0.5\]

OpenStudy (lgbasallote):

try squaring both sides...what do you get?

OpenStudy (anonymous):

is it: \[y^{2} (1-y^{2}) = 0.25\]

OpenStudy (lgbasallote):

yes. now distribute y^2 into 1 - y^2

OpenStudy (anonymous):

y^2 - y^4 = 0.25

OpenStudy (lgbasallote):

right. now subtract 0.25 from both sides

OpenStudy (anonymous):

so you get y^2 - y^4 - 0.25 = 0 and i think you have to make all the terms have the same denominatot

OpenStudy (anonymous):

\[\frac{ 4y - 4y^2 -1 }{ 4 }\]

OpenStudy (lgbasallote):

aren't you forgetting = 0?

OpenStudy (lgbasallote):

\[\frac{4y - 4y^2 - 1}4 = 0\] now...if you cross multiply.. \[4y - 4y^2 - 1 = 0\] yes?

OpenStudy (anonymous):

yeuupp now what?

OpenStudy (lgbasallote):

wait... why are you using y^2? it should be \[4y^2 - 4y^4 - 1 = 0\]

OpenStudy (lgbasallote):

now multiply by -1... \[4y^4 - 4y^2 + 1 = 0\] right?

OpenStudy (anonymous):

yes i did that and sorry typo :O Well now do we move 1 to the other side, so we subtract both sides by one?

OpenStudy (lgbasallote):

we just moved 1 to this side..why move it back?

OpenStudy (anonymous):

true, go on.

OpenStudy (lgbasallote):

now.. you perform quadratics.. let \[a = y^2\] so this becomes\[4a^2 - 4a + 1 = 0\] solve for a

OpenStudy (lgbasallote):

do you know how to?

OpenStudy (anonymous):

yes, give me a sec.

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

(2a - 1)^2 = 0

OpenStudy (lgbasallote):

right. now solve for a

OpenStudy (anonymous):

does a = 1/2 or \[a = \pm \frac{ 1 }{ 2 }\]

OpenStudy (lgbasallote):

not \(\pm\)...just +1/2

OpenStudy (lgbasallote):

so you have \[a = \frac 12\] then remember i said a = y^2 \[y^2 = \frac 12\] solve for y

OpenStudy (anonymous):

\[y = \sqrt{\frac{ 1 }{ 2 }}\]

OpenStudy (lgbasallote):

close... \[y = \pm \sqrt{\frac 12}\] ironic how you didn't use that \(\pm\) when you should have

OpenStudy (anonymous):

aha thankss man Since this is part of a system of equations, do i sub in the negative an positive values?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

substitute y = sqrt (1/2) substitute y = -sqrt(1/2)

OpenStudy (lgbasallote):

i meant part of systems of equations and substituting....oh wait i know....anyway..yes you sub

OpenStudy (anonymous):

alriight thanks man